OpenStudy (anonymous):

If the sin 30=1/2, then cos ____ = 1/2. Will give medal!

OpenStudy (anonymous):

Can someone please guide me on how to do this

zepdrix (zepdrix):

Hey Sophia! c: Sine and Cosine are co-functions. \(\Large\rm \sin(30)=\cos(60)\)

OpenStudy (anonymous):

this is from FLVS right? i would love to know the question id number

OpenStudy (anonymous):

its for a dba

OpenStudy (anonymous):

wait so the answer would be 60?

OpenStudy (anonymous):

yes, but the question is written by a moron what is dba?

OpenStudy (anonymous):

I call the teacher and let her know answers to these questions

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

tell her satellite73 says "any question that starts out with "if \(\sin(30)=\frac{1}{2}\) then \(\cos(x)=\frac{1}{2}\) " makes as much sense as saying "if \(2+2=4\) then \(\cos(x)=\frac{1}{2}\)" it was written by an idiot who has no place writing math questions

OpenStudy (anonymous):

Hah! I'll make sure to let her know! Thanks!

OpenStudy (anonymous):

really truly do it is a huge failure on the part of logic

OpenStudy (anonymous):

you don't start a question with "if blah" when "blah" is always true